# A Dataset for Large Language Model-Driven AI Accelerator Generation

### v1

@@ Abstract
In the ever-evolving landscape of Deep Neural Networks (DNN) hardware acceleration, unlocking the true potential of systolic array accelerators has long been hindered by the daunting challenges of expertise and time investment. Large Language Models (LLMs) offer a promising solution for automating code generation which is key to unlocking unprecedented efficiency and performance in various domains, including hardware descriptive code. However, the successful application of LLMs to hardware accelerator design is contingent upon the availability of specialized datasets tailored for this purpose. To bridge this gap, we introduce the Systolic Array-based Accelerator DataSet (SA-DS). SA-DS comprises of a diverse collection of spatial arrays following the standardized Berkeley’s Gemmini accelerator generator template, enabling design reuse, adaptation, and customization. SA-DS is intended to spark LLM-centred research on DNN hardware accelerator architecture. We envision that SA-DS provides a framework which will shape the course of DNN hardware acceleration research for generations to come. SA-DS is open-sourced under the permissive MIT license at this https://github.com/ACADLab/SA-DS.
@@
Systolic array design, LLM-powered hardware synthesis, accelerator architecture optimization, EDA
1 Introduction
Artificial Intelligence (AI) has shown a remarkable potential to address complex design problems ranging from software development to drug discovery. A key advantage of AI is a significant reduction of the manual effort and expertise requirements. This promising capability of AI suggests its application in hardware design, particularly for developing specialized AI accelerators needed to keep pace with the rapid evolution of Deep Neural Networks (DNNs) [1]. In hardware design for DNNs, the complexity and need for expert knowledge have been major limitations [2, 3].
Systolic array accelerators, typically obtained with specialized AI hardware generators such as Gemmini [3], have significantly advanced the processing capabilities for DNNs, providing high throughput and energy efficiency. These systems, integrated with architectures like the Rocket Chip processor [4], demonstrate the scalability and flexibility necessary for contemporary AI applications. Despite these appealing achievements, challenges such as the low-level nature, the complex programming interfaces, memory usage, and the need for extensive development times persist [5]. Moreover, systolic array accelerator generators like Gemmini [3] generally face limitations in efficiently handling diverse and irregular computational patterns beyond their optimized standard operations [6]. These limitations underscore the need for innovative solutions such as AI model-based solutions [5, 6].
At the frontier of AI advancement, Large Language Models (LLMs) [7] offer an appealing solution for alleviating the challenges in hardware accelerator design. Along this line, GPT4AIGChip [8] exemplifies using LLMs to automate the hardware design process, from conceptual design to synthesis to fabrication. However, the lack of specialized datasets of hardware accelerator design artifacts presents a strong obstacle in fully leveraging the potential of LLMs [9]. This limitation restricts usage to the vanilla LLMs without fine-tuning or in-context learning [9]; two of the most effective approaches for maximizing the LLM capabilities.
To bridge the gap, we introduce a Systolic Array Accelerator Dataset (SA-DS) to facilitate effective learning and generation of optimized designs by LLMs. Specifically, our contributions include:
(1) We create, curate, and release SA-DS, the first systolic array accelerator dataset for DNN hardware accelerator design. Each data point in SA-DS features a verbal description of an accelerator micro-architecture and a Chisel description of the design itself. These accelerator designs are obtained using the Gemmini generator [3].
(2) We demonstrate the potential of SA-DS in enabling LLM-based hardware accelerator design by showcasing its suitability for generating viable accelerator designs by providing 1-short prompts with multiple LLMs. Experimental results validate the suitability of SA-DS in providing high-quality and relevant accelerator design examples to contemporary LLMs including GPT [7], Claude [10], and Google’s Gemini [11] as compared to existing Hardware Description Language (HDL) datasets.
@@ 1 Introduction
2 Background
Static Accelerator Generation Tools: A wide array of static DNN accelerator design tools have been developed such as VTA [12], MAGNet [13], and DNNWeaver [14] to suit various applications. These tools provide many hardware architecture templates supporting vector/systolic spatial arrays, data flows, software ecosystems, and OS support. Among these tools, Gemmini [3] is a comprehensive and well-packaged open-source infrastructure tailored to design full-stack DNN accelerators. Gemmini offers a versatile hardware framework, a multi-layered software stack, and an integrated System-On-Chip (SoC) environment based on the architectural template shown in Fig. 1. Central to a design provided by Gemmini is a spatial array architecture, where the template employs a 2-D array of tiles containing Processing Elements (PE). These PE units operate in parallel, handling Multiply-Accumulate (MAC) operations efficiently. To optimize the area, power, and performance trade-offs, the size of the spatial array, input type, and function units such as non-linear functions, pooling, normalization, and dataflow can be adjusted.
@@ 2 Background
LLM for Hardware Design: LLMs are increasingly pivotal in generating HDL and High-Level Synthesis (HLS) code. Table I provides a high-level comparison of the state-of-the-art methods along this line. GitHub Copilot [19] pioneered automatic code generation, setting the stage for domain-specific applications like DAVE [20]. Expanding on these capabilities, VeriGen [15] and ChatEDA [9] advance the field by refining hardware design workflows and automating the Register-Transfer Level (RTL) to Graphic Data System version II (GDSII) process with fine-tuned LLMs. ChipGPT [16] and Autochip [18] further this evolution by integrating LLMs to generate and optimize hardware designs, with Autochip producing accurate Verilog code through simulation feedback. Chip-Chat [17] demonstrates the efficacy of interactive LLMs like ChatGPT-4 in fast-tracking design space exploration. RTLLM [21] and GPT4AIGChip [8] specifically target design process efficiency, highlighting LLMs’ capacity to manage complex design tasks and expand access to AI accelerator design, showcasing the broad potential of LLMs in hardware design. However, except for GPT4AIGChip [8], these works do not address using LLMs for AI hardware accelerator architecture design.
3 SA-DS and An Envisioned Framework
We propose a dataset for LLM-enhanced AI hardware accelerator design and envision a general framework for its applicability. Due to space limitations, we restrict our presentation to describing how it can be utilized. This framework is outlined in Fig. 2. Since an LLM’s response is determined by both the prompt and the model coefficients, the framework focuses on these two aspects. An immediate usage of SA-DS in the envisioned framework is to help to fine-tune a generic LLM for the task of hardware accelerator design (Step 2 in Fig. 2). This can be achieved by allowing its samples to partially or fully alter the model coefficients. Besides, multi-shot prompting techniques such as in-context learning can be used where SA-DS will function as the source for multi-shot examples. Also, an initial prompt conveying the user’s intent and key software and hardware specifications of the intended design can still be further engineered/optimized through the Prompt Optimizer step (Step 1). However, administering this optimization requires the development of timely and accurate evaluation techniques and metrics for the designs generated. Since SA-DS combines verbal description and systolic array design pairs, a systolic array accelerator is taken as an outcome from the LLM (step 3). Next, a third-party quality evaluation tool can be utilized to provide a quantitative evaluation of the design, verify functional correctness, and integrate the design with the full stack. (step 4).
@@ 3 SA-DS and An Envisioned Framework
In the proposed usage framework, once the LLM generates systolic array accelerators (step 3), the process moves to quality and functional evaluation (step 4). The design, often formulated in Chisel, undergoes conversion to RTL using tools like Verilator to assess functional correctness and full-stack integration, thus translating into a verifiable HDL. This step crucially embeds an automated RTL-GDSII validation process, where generated designs are assessed and flagged as Valid or Invalid based on their code sequence completeness and input-output correctness. Valid designs advance to resource validation, focusing on optimizing for Power, Performance, and Area (PPA) metrics. Conversely, designs flagged as Invalid trigger a feedback loop for error analysis and LLM retraining, facilitating iterative refinement (Steps 2 to 5) to meet the set performance criteria. Ultimately, the process culminates in Step 6, where a script generates Tool Command Language (TCL) instructions to automate RTL evaluation for HDL codes, integrating with synthesis tools to comprehensively assess and validate the PPA metrics of the hardware design.
4 SA-DS Sample Creation
@@
SA-DS uses the Gemmini generator to provide a variety of spatial array designs, making it easier for users to adapt and reuse these designs for different projects. The dataset and design tools are developed with Chisel, a programming language embedded in Scala [22], known for its clear and efficient coding style [23]. Gemmini’s configurable nature allows for significant customization, suiting various application-specific requirements, thus supporting the advancement in AI chip design [24]. This combination of a versatile template like Gemmini and a powerful design language like Chisel ensures that SA-DS can effectively meet the diverse needs of hardware design in AI applications.
@@
Algorithm 1 describes how SA-DS is created within the Chipyard framework [25], which ensures that the designs are verifiable. The process focuses on generating spatial array structures and function units from the Gemmini codebase. Specifically, the algorithm iterates through various configurations of these elements, as indicated in line 6, guided by insights from analyzing the Gemmini code. Each modification made during this process is checked for accuracy using Verilator, ensuring that each version of the design (M) is properly annotated to highlight its key features. The variables and their values used in this algorithm are carefully chosen based on extensive testing with the Gemmini template, leading to a diverse set of potential Gemmini designs, as shown in Fig. 3.
Algorithm 1 SA-DS Creation with the Gemmini Generator
1:Input: Source Code $S$
2:Output: List of Verified Modified Source Codes $M$
3:$P\leftarrow$ list of changeable variable parameters
4:$M\leftarrow$ empty list
5:function GenerateVariations($S,P$)
6:    for each $combination$ in $P$ do
7:         $S_{mod}\leftarrow S$
8:         for each $(parameter,value)$ in $combination$ do
9:             Replace $parameter$ in $S_{mod}$ with $value$
10:         end for
11:         $verified\leftarrow$ VerifyWithVerilator($S_{mod}$)
12:         if $verified$ then
13:             $M.append(S_{mod})$
14:         end if
15:    end for
16:    return $M$
17:end function
18:function VerifyWithVerilator($S_{mod}$)
19:    return Verilator verification result for $S_{mod}$
20:end function
The SA-DS, generated as detailed in Algorithm 1, offers a variety of configurations influenced by Function Unit (FU) availability and spatial array sizes, leading to a structured dataset easily navigable and applicable to diverse hardware design needs. As illustrated in Fig. 3, the dataset organizes these configurations into six main categories, each containing 1536 unique samples. These samples are enriched with dataflow variations like Output Stationary (OS), Weight Stationary (WS), and their combinations, accounting for 512 samples per dataflow type. The distribution of these parameters and their corresponding function units is systematically represented in Fig. 4, facilitating an understanding of the dataset’s comprehensive nature and the interplay between different function units in each configuration.
Figure 3: The design space parameters of the proposed SA-DS based on Gemmini[3].
Significance of the Parameters: M represents the configurations derived from Gemmini, maintaining full-stack compatibility. These configurations are crucial for defining the hardware accelerator’s micro-architectural elements based on Gemmini’s template, including scratchpad and accumulator sizes. Key parameters include:
  • Spatial Array Size: Defines the number of PEs, crucial for computational capacity.
  • DataFlow: Manages data movement among PEs, with options like OS, WS, or an automatic selection during runtime.
  • Function Units: Additional units that support DNN functionalities like ReLU and normalization.
  • Accumulation & Spatial Array Output Type: Affects computation precision, primarily supporting signed integer types, with potential expansion to floating-point and complex integer types.
These elements facilitate customization to meet specific application requirements.
@@
Figure 4: The frequency of sample sin SA-DS in terms of (a) function units in each category based on the Data flow for systolic arrays, (b) Function units available individually or in combination with the others.
5 Experimental Analysis
In this section, we evaluate the effectiveness of SA-DS in supporting the design generation process for hardware accelerators via LLMs, referencing the framework depicted in Fig. 2. Due to space limitations and given the complexity of LLM fine-tuning and prompt optimization requiring further research, our analysis is conducted conceptually. We initiate a proof-of-concept experiment that benchmarks SA-DS against a recent HLS Dataset (HLSD) from [26]. This experiment considers utilizing each dataset to supply one-shot examples for LLM prompts, aiming to enhance the generation of hardware designs from verbal descriptions. To objectively assess the impact of each dataset, we analyze the code quality derived from representative prompt-code pairs selected from SA-DS. The experiment extends across four prominent LLMs; GPT-4 [7], GPT-3.5 [27], Claude [10], and Gemini Advanced [11]. Reflecting the diversity of hardware specifications covered by SA-DS, our methodology includes randomly selecting test sets from six categories within SA-DS, ensuring each category is represented by 30 samples.
Evaluation is conducted by manual code review with an HLS and Chisel expert. Due to the manual nature of verification, a bi-state verification scheme is adopted. A pass characterizes the generation of a complete and functional code complying with the verbal description or the case where the LLM generates the most crucial portions of the code and leaves redundant lines or if the code is extended to have other unended functionalities beyond what is requested in the verbal description. Conversely, a fail refers to generating incomplete code, incorrect file headers, or incurring fatal errors of different types that render the code unfunctional. Therefore, we use Verilator as an automated design verification tool exclusively for codes marked as Pass. The results of this experiment are summarized in Table II.
TABLE II: Suitability of 1-shot examples: SA-DS vs. HLSD
LLM SA-DS HLSD
Pass Fail Pass Fail
GPT-4 135 45 72 108
Gemini Advanced 144 36 57 123
GPT-3.5 155 25 68 112
Claude 150 30 71 109
The comparison between SA-DS and HLSD datasets in generating one-shot prompts for LLMs like GPT-4, Gemini Advanced, GPT-3.5, and Claude, as shown in Table II, reveals a clear pattern. SA-DS consistently shows fewer failures and more passes across all tested LLMs with around 46% more passes on average. This suggests that SA-DS’s examples better align with the LLMs’ capabilities, leading to more effective code generation. The higher Pass rates in SA-DS imply that while not perfect, the generated code often needs fewer revisions to meet design requirements, indicating its practical value in streamlining the accelerator design process.
6 Conclusion

### v2

# SA-DS: A Dataset for Large Language Model-Driven AI Accelerator Design Generation

@@ Abstract
In the ever-evolving landscape of Deep Neural Networks (DNN) hardware acceleration, unlocking the true potential of systolic array accelerators has long been hindered by the daunting challenges of expertise and time investment. Large Language Models (LLMs) offer a promising solution for automating code generation which is key to unlocking unprecedented efficiency and performance in various domains, including hardware descriptive code. The generative power of LLMs can enable the effective utilization of preexisting designs and dedicated hardware generators. However, the successful application of LLMs to hardware accelerator design is contingent upon the availability of specialized datasets tailored for this purpose. To bridge this gap, we introduce the Systolic Array-based Accelerator DataSet (SA-DS). SA-DS comprises a diverse collection of spatial array designs following the standardized Berkeley’s Gemmini accelerator generator template, enabling design reuse, adaptation, and customization. SA-DS is intended to spark LLM-centered research on DNN hardware accelerator architecture. We envision that SA-DS provides a framework that will shape the course of DNN hardware acceleration research for generations to come. SA-DS is open-sourced under the permissive MIT license at https://github.com/ACADLab/SA-DS.
@@
Systolic array design, LLM-powered hardware synthesis, accelerator architecture
I Introduction
Artificial Intelligence (AI) has shown a remarkable potential to address complex design problems ranging from software development to drug discovery. A key advantage of AI is the significant reduction of manual effort and expertise requirements. This promising capability of AI suggests its potential in hardware design, particularly for developing specialized AI accelerators needed to keep pace with the rapid evolution of Deep Neural Networks (DNNs) [1]. In DNN hardware design, the complexity and need for expert knowledge have been major limitations [2, 3].
A wide range of static DNN accelerator design tools have been developed such as VTA [4], MAGNet [5], and DNNWeaver [6] to suit various applications. These tools provide many hardware architecture templates supporting vector/systolic spatial arrays, data flows, software ecosystems, and Operating System (OS) support. Among these tools, Gemmini [3] stands out as a comprehensive and well-packaged generator with open-source infrastructure tailored to designing full-stack DNN accelerators. Gemmini provides a versatile hardware framework, a multi-layered software stack, and an integrated System-on-Chip (SoC) environment based on the architectural template shown in Fig. 1. At the core of Gemmini’s design is a spatial array architecture, utilizing a 2-D array of tiles that contain Processing Elements (PEs). These PEs operate in parallel, efficiently handling Multiply-Accumulate (MAC) operations. Despite these appealing attributes, however, challenges such as the low-level nature, the complex programming interfaces, excessive memory usage, and the need for extensive development times persist [7]. Moreover, systolic array accelerator generators like Gemmini generally face limitations in efficiently handling diverse and irregular computational patterns beyond their optimized standard operations [8]. These limitations underscore the need for innovative solutions such as AI model-based solutions [7, 8].
@@ I Introduction
At the frontier of AI models, Large Language Models (LLMs) [9] offer an appealing solution for alleviating the challenges in hardware accelerator design. LLMs show undoubted promise in generating HDL and High-Level Synthesis (HLS) code. Table I compares notable methods in this field. GitHub Copilot [10] introduces automatic code generation, paving the way for tools like DAVE [11]. Building on these efforts, VeriGen [12] and ChatEDA [13] refine hardware design workflows, automating the RTL to GDSII process with fine-tuned LLMs. ChipGPT [14] and Autochip [15] integrate LLMs to generate and optimize hardware designs, with Autochip producing precise Verilog code through simulation feedback. Chip-Chat [16] demonstrates interactive LLMs like ChatGPT-4 in accelerating design space exploration. MEV-LLM [17] proposes multi-expert LLM architecture for Verilog code generation. Assert-O [18] is an automated framework to derive security properties from SoC documentation and optimize them. RTLLM [19] and GPT4AIGChip [20] enhance design efficiency, showcasing LLMs’ ability to manage complex design tasks and broaden access to AI accelerator design. However, most works, except GPT4AIGChip [20], do not address AI hardware accelerator architecture design challenges. Besides, the absence of specialized datasets for hardware accelerator design artifacts poses a significant barrier to fully harnessing the potential of LLMs [13]. This limitation confines their application to standard LLMs without fine-tuning or In-Context Learning (ICL) [13], which are among the most promising methods for optimizing LLM capabilities [21].
To bridge the gap, we introduce a Systolic Array Accelerator Dataset (SA-DS) to facilitate effective learning and generation of optimized designs by LLMs. Our contributions are as follows: (1) We create, curate, and release SA-DS, the first dataset of systolic array accelerators tailored for LLM-based DNN hardware accelerator design. Each data point in SA-DS includes a natural language description of the accelerator’s micro-architecture and a Chisel description, which is an HDL representation of the design. These accelerator designs are prepared using dedicated hardware generators such as the Gemmini generator. The key advancement over existing generator-based approaches is that SA-DS offers a more effective alternative by allowing LLMs to utilize pre-existing designs instead of relying solely on hardware generators. (2) We demonstrate the potential of SA-DS in facilitating LLM-based hardware accelerator design by showcasing its effectiveness in generating viable accelerator designs using single- and multi-shot learning with multiple LLMs. Experimental results validate the suitability of SA-DS in providing high-quality and relevant accelerator design examples to contemporary LLMs, including GPT-4o [22], GPT-3.5 Turbo [23], Claude 3 Opus [24], and Google’s Gemini [25], compared to existing HDL datasets.
@@
II SA-DS and An Envisioned Framework
We propose a dataset for LLM-enhanced AI hardware accelerator design and envision a framework for its applicability. Due to space limitations, we restrict our presentation to describing how it can be utilized. This framework is outlined in Fig. 2. Since an LLM’s response is determined by both the prompt and the model coefficients, the framework focuses on these two aspects. An immediate usage of SA-DS in the envisioned framework is to help fine-tune a generic LLM for the task of hardware accelerator design (Step 2 in Fig. 2). Equivalently, one may also employ ICL also known as multi-shot learning as a more computationally efficient compromise to fine-tuning [21]. Besides, multi-shot prompting techniques such as ICL can be used where SA-DS will function as the source for multi-shot examples. Also, an initial prompt conveying the user’s intent and key software and hardware specifications of the intended design can still be further engineered/optimized through the Prompt Optimizer step (Step 1). However, administering this optimization requires the development of timely and accurate evaluation techniques and metrics for the designs generated.
Since SA-DS combines verbal description and systolic array design pairs, a systolic array accelerator is taken as an outcome from the LLM (step 3). Next, a third-party quality evaluation tool can be employed to provide a quantitative evaluation of the design, verify functional correctness, and integrate the design with the full stack. (step 4).
@@ II SA-DS and An Envisioned Framework
In the proposed framework, once the LLM generates systolic array accelerators (step 3), the process moves to quality and functional evaluation (step 4). The design, often formulated in Chisel, undergoes conversion to RTL using tools like Verilator [26] to assess functional correctness and full-stack integration, thus translating into a verifiable HDL. This step crucially embeds an automated RTL-GDSII validation process, where generated designs are assessed and flagged as Valid or Invalid based on their code sequence completeness and input-output correctness. Valid designs advance to resource validation, focusing on optimizing for Power, Performance, and Area (PPA) metrics. Conversely, designs flagged as Invalid trigger a feedback loop for error analysis and LLM retraining, facilitating iterative refinement (Steps 2 to 5) to meet some preset performance criteria. Ultimately, the process culminates in Step 6, where a script generates Tool Command Language (TCL) instructions to automate RTL evaluation for HDL codes, integrating with synthesis tools to comprehensively assess and validate the PPA metrics of the hardware design.
III SA-DS Sample Creation
Figure 3: An example of one category and its design space parameters in the proposed SA-DS.
@@ III SA-DS Sample Creation
SA-DS uses the Gemmini generator to provide a variety of spatial array designs, making it easier for users to adapt and reuse these designs for different use cases. The dataset and design tools are developed with Chisel, a programming language embedded in Scala [27], known for its clear and efficient coding style [28]. Gemmini’s configurable nature allows for significant customization, meeting various application-specific requirements, thus supporting the advancement in AI chip design [29]. This combination of a versatile template like Gemmini and a powerful design language like Chisel promises that SA-DS can effectively meet the diverse needs of hardware design in AI applications.
@@
SA-DS is created within the Chipyard framework [30], which ensures that the designs are verifiable. The process focuses on generating spatial array structures and function units from the Gemmini codebase. We refer the interested reader to the dataset’s web page at https://github.com/ACADLab/SA-DS for an algorithmic description of the steps of the generation process. The variables and their values are carefully chosen based on extensive testing with Gemmini, leading to a diverse set of potential Gemmini designs, as shown in Fig. 3.
@@
Figure 4: The frequency of sample in SA-DS in terms of (a) function units in each category based on the Data flow for systolic arrays, (b) Function units available individually or in combination with the others.
SA-DS offers various configurations influenced by Function Unit (FU) availability and spatial array sizes, resulting in a structured and diverse dataset. This dataset applies to diverse hardware design needs. We carefully vary the computational and other FUs without altering the memory parameters. Due to the absence of an evaluator for obtaining design metrics, the memory parameters for any design can be either minimized or maximized based on the DNN workload. Therefore, we fixed this parameter. It is noted that when the memory parameter is fixed, Gemmini provides up to six possible design variations. Hence, we decided to exploit these six variations for each fixed memory parameter, resulting in six broad design categories. To enhance clarity, we categorized these sizes, with each category and its respective FUs depicted in Fig. 3. Each category includes six FU options. Given that these options are binary, there are $642^{6}=$ 64 possible combinations. Metrics such as input type, calculated as $64\times 8=512$, represent the enriched number of data points available in each dataflow type. Considering additional dataflow variations such as Output Stationary (OS) and Weight Stationary (WS), each category encompasses $512\times 3=1536$ data points as illustrated in Fig. 4(a). The distribution of these parameters and their corresponding FUs is illustrated in Fig. 4(b), facilitating an understanding of the dataset’s comprehensive nature and the interplay between different function units in each configuration. The overview in both dataflow and FU from Fig. 4 depicts a Gaussian distribution, which showcases the fairness of the dataset.
Figure 5: (a) The data point structure, (b) An example prompt and the output generated by the LLM.
IV Experimental Results
In this section, we evaluate the effectiveness of SA-DS in supporting the design generation process for hardware accelerators via LLMs, referencing the framework depicted in Fig. 2. Due to space limitations and the complexity of LLM fine-tuning and prompt optimization, we conduct a proof-of-concept experiment comparing SA-DS with a recent HLS Dataset (HLSD) [31] [^1] Other hardware datasets, such as [19], focus on small designs like adders and multipliers, making direct comparisons unfair and difficult.. We use each dataset to supply one-shot and multi-shot examples for LLM prompts to enhance hardware design generation from verbal descriptions. A design’s verbal description is included in an instructional prompt given to the LLM for generation, as shown in Fig. 5.
To objectively assess the impact of each dataset, we analyze the code quality derived from representative prompt-code pairs selected from SA-DS across four prominent LLMs: GPT-4o [22], GPT-3.5 [32], Claude 3 Opus [24], and Gemini Advanced [25]. Reflecting the diversity of hardware specifications covered by SA-DS, for single-shot our methodology includes randomly selecting test sets from six categories within SA-DS, ensuring each category is represented by 30 samples. For multi-shot, three data points are randomly picked and queried to generate the needed by choosing the best suitable fit from the provided examples. Evaluation is conducted through a manual code review by an HLS and Chisel expert using a bi-state verification scheme. The results are summarized in Table II. A Pass indicates the generation of complete and functional code that complies with the verbal description, or when the LLM generates the most crucial portions of the code, even if some redundant lines or extended functionalities are present. A Fail refers to incomplete code, incorrect file headers, or fatal errors rendering the code nonfunctional. For accuracy, Verilator is used as an automated design verification tool exclusively for codes marked as Pass for SA-DS. We focus on multi-shot learning rather than fine-tuning, as recent literature suggests that multi-shot learning is an efficient alternative to fine-tuning [^2] Dai et al. [21] report that ICL can be understood as implicit fine-tuning, demonstrating that ICL produces meta-gradients through forward computation similarly to how explicit fine-tuning updates model parameters through back-propagation..
TABLE II: Analysis of SA-DS vs. HLSD with single-/multi-shot prompts.
Input Prompt Single-shot Multi-shot
Datasets SA-DS HLSD SA-DS HLSD
LLMs Pass Fail Pass Fail Pass Fail Pass Fail
GPT-4o 165 15 92 88 180 0 69 111
Gemini Advanced 144 36 57 123 180 0 114 66
GPT-3.5 155 25 68 112 54 126 42 138
Claude-3.5-sonnet 150 30 71 109 180 0 57 123
The results in Table II highlight the superior performance of SA-DS over HLSD in generating hardware designs with LLMs. For single-shot prompts, SA-DS achieves significantly higher pass rates across all tested LLMs. For example, GPT-4o and Gemini Advanced show pass rates of 91.7% and 80.0% with SA-DS, compared to 51.1% and 31.6% with HLSD. Similarly, multi-shot prompts with SA-DS yield perfect pass rates (100%) for GPT-4o, Gemini Advanced, and Claude, significantly outperforming HLSD. However, we observe a considerable and comparable fail on GPT-3.5 for both SA-DS and HLSD. Our observation indicates that SA-DS examples are better aligned with the LLMs’ capabilities, leading to more effective and reliable code generation. The higher pass rates indicate that designs generated using SA-DS typically require fewer revisions, thereby streamlining the accelerator design process. This efficiency suggests that SA-DS is highly effective in improving the overall workflow. Consequently, SA-DS showcases significant practical value by enhancing both the efficiency and quality of hardware design generation when utilizing LLMs. This reduction in necessary revisions not only saves time but also minimizes resource expenditure, underscoring the dataset’s importance in AI hardware design. Furthermore, the improved quality of the initial designs can lead to faster deployment and potentially higher performance of the final products. Besides, the higher pass rates show the SA-DS potential to solve reliability issues with 100% success with the multi-shot approach. This underscores its practical value in streamlining the accelerator design process, making it a valuable tool for enhancing the efficiency and quality of hardware design generation.
V Conclusion